The decomposition of ammonia on a tungsten surface at 1100 °C NH1/2 N +3/2 H is zero order in NH3 In one experiment, when the initial concentration of NH, was 1.37-102 M, the concentration of NHj dropped to 2.22 103 M afte 2.50 103 seconds had passed Based on these data, the rate constant for the reaction is M s Submit Answer Retry Entire Group 9 more group attempts remaining
Answer #1

The integrated rate law for zero order reaction is [A]0 - [A] = kt

where [A]0 = initial concentration

[A] = final concentration

k = Zero order rate constant

t = time in seconds

Given that

[A]0 = initial concentration = 1.37 x 10-2 M

[A] = final concentration = 2.22 x 10-3 M = 0.222 x 10-2 M

k = Zero order rate constant

t = time in seconds = 2.50 x 103 seconds

Hence [A]0 - [A] = kt

\Rightarrow k = ( [A]0 - [A] ) / t = (1.37 x 10-2 M -   0.222 x 10-2 M) / 2.50 x 103 = 0.4592 x 10-5 M s-1

k = 4.592 x 10-6 M s-1 = 0.000004592 M s-1